OpenStudy (anonymous):

My AMAZING tutorial c:

OpenStudy (angela210793):

lol... :D let us decide if it is amazing or not :P

OpenStudy (this):

lets see

OpenStudy (anonymous):

Here I will show you how to add, subtract, multiply, and divide c:

OpenStudy (anonymous):

2+2=4, 2-2=0, 2x2=4, 2/2=1 c: now you may be jealous.

OpenStudy (agent47):

ha-ha-ha- ha...

OpenStudy (agent47):

ha

OpenStudy (agent47):

i'm just curious, is it allowed?

OpenStudy (this):

heheheheh lamoo

OpenStudy (agent47):

posting useless info as questions?

OpenStudy (agent47):

or answers?

OpenStudy (anonymous):

xD omg I love you *lololololol*

OpenStudy (anonymous):

@CHAD159753 - Don't use the question area for irrelevant posts like this. Also, you need to change your picture. Thank you.

OpenStudy (angela210793):

How did you do that????????